how do i solve for x?


8^x=(1/16)^(x-4)

Answer by stanbon(48569)

8^x=(1/16)^(x-4)
-------------
Rewrite as follows:
2^3x = 2^[-4(x-4)]
--------
Since the bases are the same the exponents must be equal:
3x = -4(x-4)
3x = -4x+16
7x = 16
x = 16/7
=============
Cheers,
Stan H.